OpenStudy (anonymous):

Could someone help me with this problem, it is regarding related rates: A 5 meter long ladder is leaning against the side of a house. The foot of the ladder is pulled away from the house at a rate if 0.4 m/sec. Determine how fast the top of the ladder is descending when the foot of the ladder is 3 meters from the house

OpenStudy (anonymous):

i am stuck when it comes to figuring out how to put the values into the derived equation: 2x(dx)/(dt)+2y(dy)/(dt)=0

OpenStudy (bahrom7893):

Okay imagine a ladder leaning against a wall making a triangle.. the ladder is hypothenuse, and is z, the wall is one leg, x and the floor is the other leg, y. z^2 = x^2 + y^2

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

2z z' = 2x x' + 2y y' or z z' = x x' + y y' Now z' is the change in length of z, but z does not change, it's a ladder so z' = 0

OpenStudy (bahrom7893):

z', x', y' are all derivatives with respect to t by the way, it's just that z' is easier to write than dz/dt

OpenStudy (anonymous):

where the the two's go

OpenStudy (bahrom7893):

divide everything by 2

OpenStudy (anonymous):

gotcha

OpenStudy (bahrom7893):

so now if z' = 0 then z z' = x x' + y y' 0 = x x' + y y'

OpenStudy (bahrom7893):

actually draw this.. instead of calling the wall x, call the wall y.. So draw a ladder against a wall, or simply a triangle with the ladder being z, wall is y and floor is x

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

The foot of the ladder is pulled away at the rate of 0.4m/s so x is increasing, so x' = +0.4m/s

OpenStudy (bahrom7893):

Now you need to find y' when x = 3m

OpenStudy (bahrom7893):

so: x x' + y y' = 0 3 * 0.4 + y * y' = 0 or y * y' = -1.2

OpenStudy (bahrom7893):

are you okay so far?

OpenStudy (anonymous):

yes i follow

OpenStudy (bahrom7893):

okay so now all you need to know is what y is when x = 3 to find y, do: z^2 = x^2 + y^2 z = 5 (length of the ladder) and x = 3

OpenStudy (bahrom7893):

25 = 9 + y^2 25 - 9 = y^2 y^2 = 16, y = 4

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

plug that into the previous equation: y * y' = -1.2 4y' = -1.2 y' = -1.2/4 = -0.3 m/s

OpenStudy (bahrom7893):

So the top of the ladder is DESCENDING at 0.3 m/s as you can see we got a negative value so we are right since the top is descending so the height, or y, is decreasing basically..

OpenStudy (anonymous):

this is great thanks, i appreaciate it

OpenStudy (anonymous):

do you have any tips for solving these problems ?

OpenStudy (bahrom7893):

Well always draw the graph and look at what's changing and what's not...

OpenStudy (bahrom7893):

Correct graph in these problems is like.. 30% problem solved.. Correct graph and knowing what's changing(increasing/decreasing) is about 90% problem is solved.. all you need to do after that is take what you have and plug it all in..

OpenStudy (anonymous):

thanks